# Deep learning for dynamic graphs: models and benchmarks

### v1

@@ Abstract
With the aim of fostering research in the domain of dynamic graphs, at first, we survey recent advantages in learning both temporal and spatial information, providing a comprehensive overview of the current state-of-the-art in the domain of representation learning for dynamic graphs. Secondly, we conduct a fair performance comparison among the most popular proposed approaches, leveraging rigorous model selection and assessment for all the methods, thus establishing a sound baseline for evaluating new architectures and approaches.
@@ I Introduction
Despite the progress made in recent years in the field, which is mainly related to the family of Deep Graph Networks (DGNs) [7, 8], a majority of the literature works focuses on networks that are textitstatic snapshots of a phenomenon. Indeed, this is often a limitation when considering real-world processes, both natural and synthetic, where interactions evolve over time, i.e., they are dynamic in nature. Examples of such time-evolving systems can be found in social networks, where users can develop new friendships, citation networks, which are constantly updated with new publications, and in an e-commerce, where user behaviors change over time as well as their interactions with items.
@@
The present survey has a three-fold contribution. First, we propose a coherent formalization of the domain of representation learning for dynamic graphs, unifying different definitions and formalism gathered from the literature. Secondly, we provide a survey on representation learning for dynamic graphs under our unified formalism. Finally, we provide the graph learning community with a fair performance comparison among the most popular DGNs for dynamic graphs, using a standardized and reproducible experimental environment [^1] We release openly the code at https://github.com/gravins/dynamic_graph_benchmark.. Specifically, we performed experiments with a rigorous model selection and assessment framework, in which all models were compared using the same features, same datasets and the same data splits. As a by-product of our work, we also provide the community with a selection of datasets which we put forward as good candidates for the benchmarking of future works produced by the community.
Existing surveys on the topic of deep graph learning are [13, 14]. Our novel contributions with respect to such works are: (i) a broader and more up-to-date coverage of literature; (ii) a benchmark and an empirical comparison between a broad range of methods; and (iii) a statistical overview of a selected list of datasets in both discrete and continuous time setting.
@@ II-A Definitions and notations
In many practical scenarios, nodes and edges are often enriched with additional attributes. We represent node features (also known as node representation or node embedding) as a matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d_{n}}$, where $d_{n}$ is the number of available features. The $u$-th row of $\mathbf{X}$ is denoted as $\mathbf{x}_{v}$ and represents a single node’s features. Similarly, we represent edge features as a matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$, where $d_{e}$ is the number of edge features, and we indicate edge features vectors as $\mathbf{e}_{uv}$. Finally, we denote the neighborhood (or adjacency set) of a node $u\in\mathcal{V}$ as the set $\mathcal{N}_{u}=\{v\in\mathcal{V}\mid\{u,v\}\in\mathcal{E}\}$. A visual representation of a directed and undirected graph and the neighborhood of a node is shown in Figure 1.
@@ II-B1 Spectral convolution
$\mathbf{g}_{\theta}*\mathbf{x}=\mathbf{U}\mathbf{g}_{\theta}\mathbf{U}^{T}\mathbf{x}$ (1)
@@
$\mathbf{x}^{\ell+1}_{u}=\mathbf{\Theta}_{0}\mathbf{x}_{u}^{\ell}+\mathbf{\Theta}_{1}\sum_{v\in\mathcal{N}_{u}}\frac{1}{\sqrt{\mathbf{deg}(v)\,\mathbf{deg}(u)}}\mathbf{x}^{\ell}_{v},$ (2)
@@ II-B2 Spatial convolution
$\mathbf{x}_{u}^{\ell+1}=\phi_{U}(\mathbf{x}_{u}^{\ell},\bigoplus_{v\in\mathcal{N}_{u}}\phi_{M}(\mathbf{x}_{u}^{\ell},\mathbf{x}_{v}^{\ell},\mathbf{e}_{uv}))$ (3)
@@
$\mathbf{x}_{u}^{\ell+1}=\sigma\left(\sum_{v\in\mathcal{N}_{u}}\alpha_{uv}\mathbf{\Theta}\mathbf{x}_{v}^{\ell}\right)$ (4)
@@
When graphs are large and dense, i.e., $|\mathcal{E}|=|\mathcal{V}|^{2}$, it can be impractical to perform the convolution over the node’s neighborhood. Neighborhood sampling has been proposed as a possible strategy to overcome this limitation, i.e. by using only a random subset of neighbors to update node representation. GraphSAGE [21] exploits this strategy to improve efficiency and scale to large graphs. GraphSAGE updates the representation of a node $u$ by fixing the subset of nodes treated as neighbors, and by leveraging aggregation and concatenation operations:
$\mathbf{x}_{u}^{\ell+1}=\sigma(\mathbf{\Theta}\cdot[\mathbf{x}_{u}^{\ell}\,||\,\square(\{\mathbf{x}_{v}^{\ell}\}_{v\in\mathcal{N}_{S}(u)}])$ (5)
where $\mathcal{N}_{S}:\mathcal{V}\rightarrow\mathcal{V}$ is the function that computes the fixed subset of neighbors for a node $u$, and $\square$ is am aggregation function. Differently, ClusterGCN [22] samples a block of nodes identified by a graph clustering algorithm to restrict the neighborhood dimension.
@@
$\mathbf{x}_{u}^{\ell+1}=\mathrm{MLP}\left((1+\gamma)\mathbf{x}_{u}^{\ell}+\sum_{v\in\mathcal{N}_{u}}\mathbf{x}_{v}^{\ell}\right)$ (6)
@@ III Representation learning for dynamic Graphs: Notation and Taxonomy
A dynamic graph is a tuple $\mathcal{G}(t)=(\mathcal{V}(t),\mathcal{E}(t),\mathbf{X}(t),\mathbf{E}(t))$, defined for $t\geq 0$. Differently from static graphs, all elements in the tuple are functions of time $t$. Thus, $\mathcal{V}(t)$ provides the set of nodes which are present in the graph at time $t$, and $\mathcal{E}(t)\subseteq\{\{u,v\}\,|\,u,v\in\mathcal{V}(t)\}$ defines the links between them. Analogously, $\mathbf{X}(t)$ and $\mathbf{E}(t)$ define node states and edge attributes at time $t$. Although, $\mathcal{V}(t)$ can theoretically change over time, in practice it is often considered fixed for the ease of computation, which means that all the nodes that will appear in the dynamic graph are known in advance. Hence, $\mathcal{V}(t)=\mathcal{V}$ for $t\geq 0$.
The way we observe a system of interacting entities plays a crucial role in the definition of the corresponding dynamic graph. We can distinguish between two distinct types: discrete-time dynamic graphs and continuous-time dynamic graphs. Each of these representations gives rise to diverse architectures and learning approaches.
A discrete-time dynamic graph (D-TDG) models an evolving system that is fully observed at different timestamps. For such a reason, a D-TDG, $\mathcal{G}=\{\mathcal{G}_{t}\,|\,t\in[t_{0},t_{n}]\}$, consists in a sequence of static graphs (known as snapshots). Each snapshot, $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t},\mathbf{E}_{t})$, provides a picture of the graph’s state at a particular time $t$. Each snapshot maintains the notations and definitions outlined for static graphs (see Section II-A). We note that if the set of nodes and edges are fixed over time (i.e., $\mathcal{G}_{t}=(\mathcal{V},\mathcal{E},\mathbf{X}_{t},\mathbf{E}_{t})$), then the dynamic graph is often referred to as spatio-temporal graph.
@@
A continuous-time dynamic graph (C-TDG) is a more general formulation of a dynamic graph. Indeed, it models systems that are not fully observed over time. In facts, only new events in the system are observed. Therefore, a C-TDG is a stream of events (also known as observations) $\mathcal{G}=\{o_{t}\,|\,t\in[t_{0},t_{n}]\}$. An event, $o_{t}=(t,\,EventType,\,\{u\}_{u\in\mathcal{V}(t)})$, is a tuple containing the timestamp, the event type, and the nodes involved. Since C-TDGs are captured with fine-grain observations, they typically have irregular timestamps. Without loss of generality, we can identify three types of event:
  • node-wise, i.e., a node is created or its features are updated;
  • interaction, i.e., a temporal edge is created;
  • deletion, i.e., a node or an edge is deleted.
Generally, the temporal neighborhood of a node $u$ at time $t$, consists of all the historical neighbors of $u$, prior to current time $t$, i.e., $\mathcal{N}^{t}_{u}=\{v\in\mathcal{V}(t)\,|\,(t_{i},\,interaction,\,u,v)\in\mathcal{G}\wedge t_{i}<t\}$.
@@ IV Learning with Discrete-Time Dynamic Graphs
Given the sequential structure of D-TDGs, a natural choice for many methods has been to extend Recurrent Neural Networks [33] to graph data. Indeed, most of the models presented in the literature can be summarized as a combination of static DGNs and RNNs. In particular, some approaches adopt a stacked architecture, where DGNs and RNNs are used sequentially, enabling to separately model spatial and temporal dynamics. Other approaches integrate the DGN inside the RNN, allowing to jointly capture the temporal evolution and the spatial dependencies in the graph. In the following, we review state-of-the-art approaches for both spatio-temporal graphs and more general D-TDGs.
@@ IV-A1 Stacked architectures
Seo et al. [34] proposed Graph Convolutional Recurrent Network (GCRN), one of the earliest deep learning models able to learn sequences of spatio-temporal graphs. The authors proposed to stack a Chebyshev spectral convolution [18] (Equation 2 shows the first-order approximation of this convolution) for graph embedding computation and a Peephole-LSTM [35, 36] for sequence learning:
@@
$\begin{split}\mathbf{X}^{\prime}_{t}&=\text{{dgn}}(\mathbf{X}_{t},\mathbf{\Theta})\\ \mathbf{H}_{t}&=\text{{rnn}}(\mathbf{X}^{\prime}_{t})\end{split}$ (8)
Li et al. [37] implement Equation 8 by leveraging the same spectral convolution as GCRN and a Gated Recurrent Unit (GRU) [38] as RNN. Differently, Zhao et al. [9] employed the first-order approximation of the Chebyshev polynomials, which lead to the usage of a GCN to learn spatial features, and a GRU [38] to extract temporal features. This results in:
$\begin{split}\mathbf{X}^{\prime}_{t}&=\text{{gcn}}(\mathbf{X}_{t})\\ \mathbf{H}_{t}&=\text{{gru}}(\mathbf{X}^{\prime}_{t}).\end{split}$ (9)
A3TGCN [39] extends the implementation of Zhao et al. [9] with an attention mechanism to re-weight the influence of historical node states, with the aim of capturing more global information about the system.
@@ IV-A2 Integrated architectures
$\hat{h}=\sigma(\text{Cheb}(\mathbf{X}_{t},\mathbf{\Theta}_{x})+\text{Cheb}(\mathbf{H}_{t-1},\mathbf{\Theta}_{h})+\theta_{c}\odot c_{t-1}),$ (10)
@@
$\mathbf{H}_{t}=\mathrm{Conv}^{\mathcal{T}}_{2}\left(\,\mathrm{ReLU}(\,\mathrm{Conv}^{\mathcal{G}}(\,\mathrm{Conv}^{\mathcal{T}}_{1}(\mathbf{X}_{t})\,))\right).$ (11)
@@ IV-B1 Integrated architectures
$\hat{h}_{t}=\sigma(\mathbf{\Theta}_{h}\mathbf{A}_{t}+\text{{gcn}}_{h}(\mathbf{H}_{t-1},\mathcal{E}_{t-1}))$ (12)
@@
A similar strategy, which integrates topological changes into the computation, has also been employed by Li et al. [43]. Indeed, the authors proposed the so-called LRGCN that embed a Relational-GCN [44] into a LSTM model. Different from GC-LSTM, LRGCN exploits the directionality of the edges in accordance with node features rather than the only stream of adjacency matrices, with the aim of effective modeling of the temporal dynamics. In LRGCN the input, forget, and output gates are computed as the result of the R-GCN model over the input node representations and the node embeddings computed at the previous step. The authors distinguish between four edge types to produce more informed latent representation: intra-incoming, intra-outgoing, inter-incoming, inter-outgoing. An inter-time relation corresponds to an arc $\{u,v\}$ present at time $t-1$, while an intra-time relation is an arc $\{u,v\}$ present at time $t$. To employ LRGCN in a path classification task, the authors extend their model with a self-attentive path embedding (SAPE). Given a path $P\in\mathbb{R}^{m\times d_{o}}$, where $m$ is the path length and $d_{o}$ is the output dimension, SAPE first applies LSTM to capture node dependency along the path sequence, i.e., $\Gamma=\mathrm{LSTM}(P)\,\in\mathbb{R}^{m\times d_{new}}$. Then, SAPE uses the self-attentive mechanism to learn node importance and generate size-invariant representation of the path, i.e.,
$S=\mathrm{softmax}(\mathrm{MLP}(\mathrm{tanh}(\mathrm{MLP}(\Gamma))))\,\in\mathbb{R}^{r\times m}$ (13)

### v2

@@ Abstract
With the aim of fostering research in the domain of dynamic graphs, at first, we survey recent advantages in learning both temporal and spatial information, providing a comprehensive overview of the current state-of-the-art in the domain of representation learning for dynamic graphs. Secondly, we conduct a fair performance comparison among the most popular proposed approaches on node and edge-level tasks, leveraging rigorous model selection and assessment for all the methods, thus establishing a sound baseline for evaluating new architectures and approaches.
@@ I Introduction
Despite the progress made in recent years in the field, which is mainly related to the family of Deep Graph Networks (DGNs) [7, 8], a majority of the literature works focuses on networks that are static snapshots of a phenomenon. Indeed, this is often a limitation when considering real-world processes, both natural and synthetic, where interactions evolve over time, i.e., they are dynamic in nature. Examples of such time-evolving systems can be found in social networks, where users can develop new friendships, citation networks, which are constantly updated with new publications, and in an e-commerce, where user behaviors change over time as well as their interactions with items.
@@
The present survey has a three-fold contribution. First, we propose a coherent formalization of the domain of representation learning for dynamic graphs, unifying different definitions and formalism gathered from the literature. Secondly, we provide a survey on representation learning for dynamic graphs under our unified formalism (the taxonomy behind our surveying methodology is depicted in Figure 1). Finally, we provide the graph learning community with a fair performance comparison among the most popular DGNs for dynamic graphs, using a standardized and reproducible experimental environment [^1] We release openly the code at https://github.com/gravins/dynamic_graph_benchmark.. Specifically, we performed experiments with a rigorous model selection and assessment framework, in which all models were compared using the same features, same datasets and the same data splits. As a by-product of our work, we also provide the community with a selection of datasets which we put forward as good candidates for the benchmarking of future works produced by the community.
Existing surveys on the topic of deep graph learning are [13, 14]. Our novel contributions with respect to such works are: (i) a broader and more up-to-date coverage of literature; (ii) a benchmark and an empirical comparison between a broad range of methods; and (iii) the identification of a curated list of datasets in both discrete and continuous time setting.
@@
Deep Graph NetworksStatic $*{}^{*}$ptSpectral(Sec II.B.1)ptSpatial(Sec $*{}^{*}$or a deeper analysis we refer to [7, 8]ptRandom walks(Sec II.B.3)Dynamic graphsD-TDGsSpatio-temporalptStacked(Sec IV.A.1)ptIntegrated(Sec IV.A.2)Generic D-TDGsptIntegrated(Sec IV.B.1)ptStacked(Sec IV.B.2)ptMeta(Sec IV.B.3)ptAutoencoder(Sec IV.B.4)ptRandom walks(Sec IV.B.5)C-TDGsptIntegrated(Sec V.1)ptStacked(Sec V.2)ptRandom walks(Sec V.3)ptHybrid(Sec V.4)
Figure 1: Taxonomy employed to structure our survey of TDG models.
@@ II-A Definitions and notations
In many practical scenarios, nodes and edges are often enriched with additional attributes. We represent node features (also known as node representation or node embedding) as a matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d_{n}}$, where $d_{n}$ is the number of available features. The $u$-th row of $\mathbf{X}$ is denoted as $\mathbf{x}_{v}$ and represents a single node’s features. Similarly, we represent edge features as a matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$, where $d_{e}$ is the number of edge features, and we indicate edge features vectors as $\mathbf{e}_{uv}$. Finally, we denote the neighborhood (or adjacency set) of a node $u\in\mathcal{V}$ as the set $\mathcal{N}_{u}=\{v\in\mathcal{V}\mid(v,u)\in\mathcal{E}\}$. A visual representation of a directed and undirected graph and the neighborhood of a node is shown in Figure 2.
@@ II-B1 Spectral convolution
$\small\mathbf{g}_{\theta}*\mathbf{x}=\mathbf{U}\mathbf{g}_{\theta}\mathbf{U}^{T}\mathbf{x}$ (1)
@@
$\small\mathbf{x}^{\ell+1}_{u}=\mathbf{\Theta}_{0}\mathbf{x}_{u}^{\ell}+\mathbf{\Theta}_{1}\sum_{v\in\mathcal{N}_{u}}\frac{1}{\sqrt{\mathbf{deg}(v)\,\mathbf{deg}(u)}}\mathbf{x}^{\ell}_{v},$ (2)
@@ II-B2 Spatial convolution
$\small\mathbf{x}_{u}^{\ell+1}=\phi_{U}(\mathbf{x}_{u}^{\ell},\bigoplus_{v\in\mathcal{N}_{u}}\phi_{M}(\mathbf{x}_{u}^{\ell},\mathbf{x}_{v}^{\ell},\mathbf{e}_{uv}))$ (3)
@@
$\small\mathbf{x}_{u}^{\ell+1}=\sigma\left(\sum_{v\in\mathcal{N}_{u}}\alpha_{uv}\mathbf{\Theta}\mathbf{x}_{v}^{\ell}\right)$ (4)
@@
When graphs are large and dense, i.e., $|\mathcal{E}|$close to $|\mathcal{V}|^{2}$, it can be impractical to perform the convolution over the node’s neighborhood. Neighborhood sampling has been proposed as a possible strategy to overcome this limitation, i.e. by using only a random subset of neighbors to update node representation. GraphSAGE [21] exploits this strategy to improve efficiency and scale to large graphs. GraphSAGE updates the representation of a node $u$ by fixing the subset of nodes treated as neighbors, and by leveraging aggregation and concatenation operations:
$\small\mathbf{x}_{u}^{\ell+1}=\sigma\Biggl{(}\mathbf{\Theta}\cdot\Bigl{[}\mathbf{x}_{u}^{\ell}\,\big{\|}\,\bigoplus_{v\in\mathcal{N}_{S}(u)}\mathbf{x}_{v}^{\ell}\Bigr{]}\Biggr{)}$ (5)
where $\mathcal{N}_{S}:\mathcal{V}\rightarrow\mathcal{V}$ is the function that computes the fixed subset of neighbors for a node $u$. Differently, ClusterGCN [22] samples a block of nodes identified by a graph clustering algorithm to restrict the neighborhood dimension.
@@
$\small\mathbf{x}_{u}^{\ell+1}=\mathrm{MLP}\left((1+\gamma)\mathbf{x}_{u}^{\ell}+\sum_{v\in\mathcal{N}_{u}}\mathbf{x}_{v}^{\ell}\right)$ (6)
@@ III Representation learning for dynamic Graphs: Notation and Taxonomy
A dynamic graph (also referred to as temporal graph) is a tuple $\mathcal{G}(t)=(\mathcal{V}(t),\mathcal{E}(t),\mathbf{X}(t),\mathbf{E}(t))$, defined for $t\geq 0$. Differently from static graphs, all elements in the tuple are functions of time $t$. Thus, $\mathcal{V}(t)$ provides the set of nodes which are present in the graph at time $t$, and $\mathcal{E}(t)\subseteq\{\{u,v\}\,|\,u,v\in\mathcal{V}(t)\}$ defines the links between them. Analogously, $\mathbf{X}(t)$ and $\mathbf{E}(t)$ define node states and edge attributes at time $t$. Although, $\mathcal{V}(t)$ can theoretically change over time, in practice it is often considered fixed for the ease of computation, which means that all the nodes that will appear in the dynamic graph are known in advance. Hence, $\mathcal{V}(t)=\mathcal{V}$ for $t\geq 0$.
The way we observe a system of interacting entities plays a crucial role in the definition of the corresponding dynamic graph. We can distinguish between two distinct types: discrete-time dynamic graphs and continuous-time dynamic graphs [13]. Each of these representations gives rise to diverse architectures and learning approaches.
A discrete-time dynamic graph (D-TDG) models an evolving system that is fully observed at different timestamps. For such a reason, a D-TDG, $\mathcal{G}=\{\mathcal{G}_{t}\,|\,t\in[t_{0},t_{n}]\}$, consists in a sequence of static graphs (known as snapshots). Each snapshot, $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t},\mathbf{E}_{t})$, provides a picture of the whole graph’s state at a particular time $t$. Each snapshot maintains the notations and definitions outlined for static graphs (see Section II-A).
We note that if the set of nodes and edges are fixed over time (i.e., $\mathcal{G}_{t}=(\mathcal{V},\mathcal{E},\mathbf{X}_{t},\mathbf{E}_{t})$), then the dynamic graph is often referred to as spatio-temporal graph.
@@
A continuous-time dynamic graph (C-TDG) is a more general formulation of a dynamic graph. It models systems that are not fully observed over time. In facts, only new events in the system are observed. Therefore, a C-TDG is a stream of events (observations) $\mathcal{G}=\{o_{t}\,|\,t\in[t_{0},t_{n}]\}$. An event, $o_{t}=(t,\,EventType,\,\{u\}_{u\in\mathcal{V}(t)})$, is a tuple containing the timestamp, the event type, and the nodes involved. Such events have a clear interpretation as graph edits [33], which have also been formulated for dynamic graphs [34]. We consider six edit types: node deletions, node insertions, node replacements (i.e. features change), edge deletions, edge insertions, and edge replacements (i.e., edge features change). Without loss of generality, we can group these events into three main categories: node-wise events, when a node is created or its features are updated; interaction events, i.e., a temporal edge is created; deletion events, i.e., node/edge is deleted.
Despite the underlying discrete nature of events, the granularity of the observations is refined to the extent that the dynamic graph is considered as a continuous flow, allowing for events to happen at any moment (i.e., characterized by irregular timestamps). This is in contrast to discrete-time dynamic graphs, where changes to the graph are typically aggregated.
Generally, the temporal neighborhood of a node $u$ at time $t$, consists of all the historical neighbors of $u$, prior to current time $t$, i.e., $\mathcal{N}^{t}_{u}=\{v\in\mathcal{V}(t)\,|\,(t_{i},\,interaction,\,v,u)\in\mathcal{G}\wedge t_{i}<t\}$.
@@ IV Learning with Discrete-Time Dynamic Graphs
Given the sequential structure of D-TDGs, a natural choice for many methods has been to extend Recurrent Neural Networks [35] to graph data. Indeed, most of the models presented in the literature can be summarized as a combination of static DGNs and RNNs. In particular, some approaches adopt a stacked architecture, where DGNs and RNNs are used sequentially, enabling to separately model spatial and temporal dynamics. Other approaches integrate the DGN inside the RNN, allowing to jointly capture the temporal evolution and the spatial dependencies in the graph. Thus, the primary distinction between static and dynamic approaches lies in their architectures. Static approaches predominantly utilize feedforward or recurrent architectures. Both used for exploring and learning the inherent static graph structure. Differently, dynamic approaches are characterized by recurrent architectures for learning and capturing temporal and spatial dependencies within the evolving graph, which mirrors the increased complexity inherent in addressing dynamic graphs. In the following, we review state-of-the-art approaches for both spatio-temporal graphs and more general D-TDGs.
@@ IV-A1 Stacked architectures
Seo et al. [36] proposed Graph Convolutional Recurrent Network (GCRN), one of the earliest deep learning models able to learn spatio-temporal graphs. The authors proposed to stack a Chebyshev spectral convolution [18] (Equation 2 shows the first-order approximation of this convolution) for graph embedding computation and a Peephole-LSTM [37, 38] for sequence learning:
@@
$\small\begin{split}\mathbf{X}^{\prime}_{t}&=\text{{dgn}}(\mathbf{X}_{t},\mathbf{\Theta})\\ \mathbf{H}_{t}&=\text{{rnn}}(\mathbf{X}^{\prime}_{t})\end{split}$ (8)
Li et al. [39] implement Equation 8 by leveraging the same spectral convolution as GCRN and a Gated Recurrent Unit (GRU) [40] as RNN. Differently, Zhao et al. [9] employed the first-order approximation of the Chebyshev polynomials, which lead to the usage of a GCN to learn spatial features, and a GRU to extract temporal features. A3TGCN [41] extends the implementation of Zhao et al. [9] with an attention mechanism to re-weight the influence of historical node states, with the aim of capturing more global information about the system.
@@ IV-A2 Integrated architectures
$\small\hat{h}=\sigma(\text{Cheb}(\mathbf{X}_{t},\mathbf{\Theta}_{x})+\text{Cheb}(\mathbf{H}_{t-1},\mathbf{\Theta}_{h})+\theta_{c}\odot c_{t-1}),$ (9)
@@
$\small\mathbf{H}_{t}=\mathrm{Conv}^{\mathcal{T}}_{2}\left(\,\mathrm{ReLU}(\,\mathrm{Conv}^{\mathcal{G}}(\,\mathrm{Conv}^{\mathcal{T}}_{1}(\mathbf{X}_{t})\,))\right).$ (10)
Guo et al. [44] extend the spatio-temporal blocks with an attention mechanism on both spatial and temporal dimensions to better capture the spatial-temporal dynamic of the graph.
@@ IV-B1 Integrated architectures
$\small\hat{h}_{t}=\sigma(\mathbf{\Theta}_{h}\mathbf{A}_{t}+\text{{gcn}}_{h}(\mathbf{H}_{t-1},\mathcal{E}_{t-1}))$ (11)
@@
A similar strategy, which integrates topological changes into the computation, has also been employed by Li et al. [46]. Indeed, the authors proposed the so-called LRGCN that embed a Relational-GCN [47] into a LSTM model. Different from GC-LSTM, LRGCN exploits the directionality of the edges in accordance with node features rather than the only stream of adjacency matrices, with the aim of effective modeling of the temporal dynamics. In LRGCN the input, forget, and output gates are computed as the result of the R-GCN model over the input node representations and the node embeddings computed at the previous step. The authors distinguish between four edge types to produce more informed latent representation: intra-incoming, intra-outgoing, inter-incoming, inter-outgoing. An inter-time relation corresponds to an arc $(u,v)$ present at the previous time $t-1$, while an intra-time relation is an arc $(u,v)$ present at current time $t$. To employ LRGCN in a path classification task, the authors extend their model with a self-attentive path embedding (SAPE). Given the representations of $m$ nodes on a path, $P\in\mathbb{R}^{m\times d_{o}}$ with $d_{o}$ is the output dimension of LRGCN, SAPE first applies LSTM to capture node dependency along the path sequence, i.e., $\Gamma=\mathrm{LSTM}(P)\,\in\mathbb{R}^{m\times d_{new}}$. Then, SAPE uses the self-attentive mechanism to learn node importance and generate size-invariant representation of the path, i.e.,
$\small S=\mathrm{softmax}(\mathrm{MLP}(\mathrm{tanh}(\mathrm{MLP}(\Gamma))))\,\in\mathbb{R}^{r\times m}$ (12)
@@
With the aim of speeding up the dynamic graph processing, Micheli and Tortorella [48] propose DynGESN, an extension of the Graph Echo State Network [49] to the temporal domain. Specifically, DynGESN updates the embedding for a node $u$ at time $t$ as
$\small\mathbf{h}_{t}^{u}=(1-\gamma)\mathbf{h}_{t-1}^{u}+\gamma\text{tanh}\left(\mathbf{\Theta}_{i}\mathbf{h}^{u}_{t}+\sum_{v\in\mathcal{N}^{t}_{u}}\mathbf{\Theta}_{r}\mathbf{x}^{v}_{t-1}\right),$ (13)